# MCGMark: An Encodable and Robust Online Watermark for LLM-Generated Malicious Code

### v1

@@ 5 Experiments
For RQ1, we evaluate the watermark embedding success rate and detection success rate of MCGMark across 406 prompts included in MCGTest. For RQ2, we conduct a user study to assess the impact of watermark embedding on the quality of generated code. For RQ3, we design watermark attack patterns based on adversarial scenarios and evaluate the robustness of the watermarks. For RQ4, we primarily examine the influence of key parameters on MCGMark’s embedding success rate.
5.1 Experiment Setup
Evaluation Dataset: We use the MCGTest to validate the watermarking scheme. This dataset consists of $406$ malicious code prompts, including real instances generated by LLM and malicious code from high-quality open-source repositories.
LLM Selection: We test our watermarking strategy on the widely known Code LLM, DeepSeek-Coder-6.7b-instruct. This is because this LLM is open-source, allowing us to access its vocabulary (including $32,022$ tokens) and adapt MCGMark accordingly. Additionally, it has demonstrated high performance with relatively small parameters in multiple benchmarks [5].
@@ 5.1 Experiment Setup
Parameter: We set the number of LLM maximum output token, $L$, to $400$. Additionally, we set $\lceil\frac{|D|}{|V|}\rceil=0.5$. The total length of the watermark, $X$, is set to $24$ bits. And the watermark information is randomly generated. $Q3$ is set to $0.75$, and $Q1$ is set to $0.25$, which are commonly used settings for calculating outliers with quartiles [57].
@@
Environment: All experiments are conduct on a workstation with $128$ CPU cores and $8$$\times$ NVIDIA A800 (80G) GPUs.
Comparison Tools: We use WLLM, the state-of-the-art LLM watermarking technique [25], as our baseline for comparison. WLLM is an open-source LLM online watermarking techniques, aligning closely with the objectives of MCGMark. Most other post-processing watermarking techniques are less suitable for code tasks and tend to impose fixed patterns on the code, making it more detectable and prone to tampering [14]. Therefore, we choose WLLM as a comparison benchmark. We adhere to the default parameter settings as specified in the paper, with $\gamma$ set to $0.25$ and $\delta$ set to $2$. To ensure fairness, the number of tokens generated by WLLM is also limited to $400$, and the same hash value as used in MCGMark is applied.
5.2 RQ1: Effectiveness of MCGMark
To answer RQ1, we evaluate the embedding success rate and detection success rate of MCGMark under a smaller token limit using the MCGTest dataset.
@@ 5.2 RQ1: Effectiveness of MCGMark
5.2.1 Watermark Embedding Success Rate
In the MCGTest dataset, MCGMark successfully embed watermarks in $361$ out of $406$ prompts, achieving an embedding success rate of approximately $88.9\%$. In contrast, WLLM successfully embed watermarks in $342$ prompts, with an embedding success rate of around $84.2\%$. MCGMark’s embedding success rate is slightly higher than WLLM’s, MCGMark’s watermarks can trace user information, while WLLM’s can only identify whether the watermark is generated by an LLM, failing to trace the source of malicious code.
@@ 5.2.1 Watermark Embedding Success Rate
We subsequently analyze the $45$ instances where MCGMark failed to embed watermarks. We find that $17$ tasks failed due to the generated code being too short to embed the watermark successfully. Additionally, $25$ tasks failed because the generated code triggered numerous traits, resulting in a high number of assignment statements and comments, which hindered watermark embedding. Another three prompts reject the malicious code generation request, resulting in empty content. Higher folding watermark encoding could potentially solve the issue of shortcode generation failures, though this is not the focus of this paper. The failures related to code structure primarily stemmed from our restriction of LLMs to generate a maximum of $400$ tokens. In practical applications, LLMs typically generate 20484096204840962048–40962048 – 4096 tokens or even more (e.g., DeepSeek-Coder-6.7b can generate up to 64K [5]). We retest the $25$ tasks that fail due to code structure with a maximum length setting of $2048$ tokens and successfully embed watermarks in $21$ of them, achieving an embedding success rate of $84\%$. Therefore, under conditions where token numbers are unrestricted or have a sufficiently high upper limit, MCGMark’s watermark embedding success rate could reach approximately $94.1\%$ across $406$ tasks.
5.2.2 False Rate Analysis
@@
Out of the $361$ tasks where MCGMark successfully embed watermarks, $353$ watermarks are detected successfully, resulting in a detection success rate of approximately $97.8\%$. In contrast, WLLM does not support false positive rate checking. The eight instances where MCGMark failed to detect watermarks can be attributed to discrepancies in token splitting by the tokenizer. This issue leads to errors when verifying tokens against the vocabulary, returning incorrect results. This limitation is inherent to SSP, and we cannot improve the detection success rate by modifying MCGMark.
Answer to RQ 1: Under the $400$ token limit, MCGMark achieves a watermark embedding success rate of approximately $88.9\%$, with a detection success rate of around $97.8\%$. This demonstrates the effectiveness of MCGMark.
5.3 RQ2: Impact on LLM Code Generation
@@
To answer RQ 2, we conduct a user study to evaluate the impact of MCGMark on the quality of LLM-generated. We primarily assess the code generation quality of the LLM to explore the impact of MCGMark on the code generation quality. In this part, we randomly select $50$ tasks from the $346$ successfully embedded tasks of MCGTest. Furthermore, we obtain $100$ code segments, with $50$ generated by the model using our watermarking strategy and $50$ without. We invite $10$ developers with at least $4$ years of development experience (excluding co-authors), including $6$ Ph.D. students, $2$ undergraduate students, and $2$ software engineers specializing in computer-related fields, to participate in our evaluation. We randomly shuffle the order of the $50$ code pairs and further shuffle the code order within each pair. We ask the developers to identify the code they believe contains a watermark for each code pair. We collect a total of $1000$ valid responses and organize the accuracy of these $1000$ responses, as shown in Table LABEL:table:rq2dis.
confidence interval $\approx\frac{\sum\text{Results}}{\text{NO. of participants}}\pm(t\times\frac{\text{Standard deviation}}{\sqrt{\text{NO. of participants}}})$ (3)
@@
The recognition accuracy of $10$ participants for $500$ pairs of watermark/unwatermark code, totaling $1000$ segments, is $47.8\%$, which is close to random sampling. Moreover, the independent recognition rates of the $10$ participants, excluding the highest value (participant 5) and the lowest value (participant 6), fluctuate between $40\%$ to $50\%$. Furthermore, all these rates fall within the $95\%$ confidence interval (with Equ.(3), which is $[21.972,27.828]$). So, we can conclude that experienced practitioners with long-term development experience cannot correctly distinguish between watermark and unwatermark codes. This also demonstrates the stealthiness of our watermark and further confirms that the impact of MCGMark on code quality can be considered negligible.
@@
Answer to RQ 2: MCGMark ensures the quality of the LLM’s code generation while embedding watermarks, without significantly affecting the model’s normal use.
TABLE III: Types and descriptions of attacks.
@@
(2) modify/remove inputs and outputs Modifying or removing output and input content.
(3) modify/remove comments Modifying or removing single-/multi-line comments.
@@
5.4 RQ3: Resistance to Tampering
In this RQ, our main focus is on the robustness of the watermark in Section 4.4. Based on the literature [77], we primarily consider the following two types of attacks, as shown in Table III. These $8$ attack types cover the majority of modifications that typically employ against code watermarks [75, 70]. We also evaluate the performance of $50$ successfully watermarked codes under $8$ attacks, particularly whether these attacks affect the embedded watermark elements. For each attack, we conduct three attack instances. We carry out a total of $1200$ attacks.
In $1200$ attacks, we achieve a complete defense against Attacks 3, 6, 7, and 8. However, Attacks 4 and 1 failed $13$ times and $9$ times, respectively, resulting in defense success rates of $91.3\%$ and $94\%$. Upon carefully examining the failed instances, we identify a flaw in our token matching strategy during the implementation of Pattern 1-7. We rely on a generic regular expression matching approach, which proves inadequate for matching the tokens generated by LLMs due to their deviation from human language rules. The same phenomenon also affects the defense against Attack 2, resulting in an accuracy rate of $93.3\%$. Fortunately, this issue can be resolved by designing a more powerful matching scheme or SSP adapting the token vocabulary for the watermark.
@@ 5.4 RQ3: Resistance to Tampering
Additionally, it is worth emphasizing that malicious developers may attempt to modify the code generated by other LLMs. However, as long as the generated LLM incorporates our watermark, the watermark information cannot be removed from the code.
@@ 5.5 RQ4: Impact of Settings
We explore the impact of three hyperparameters on the success rate of watermark embedding. We randomly select $50$ prompts from the $406$ tasks. Under the premise of unchanged settings elsewhere, we first investigate the impact of $\lceil\frac{|D|}{|V|}\rceil$ on the success rate of watermark embedding, where $\lceil\frac{|D|}{|V|}\rceil$ represents the ratio of vocabulary $\mathbb{A}$ or $\mathbb{B}$ in the total vocabulary during vocabulary partitioning. We fix the maximum output length of the LLM at $400$ and set $\lceil\frac{|D|}{|V|}\rceil$ to $[0.25,0.375,0.5,0.625,0.75]$, respectively. The setting of $\lceil\frac{|D|}{|V|}\rceil$ is based on the observed values of the literature [25]. Our results are shown in Fig. LABEL:fig:rq1. (a). The results show that the value of $\lceil\frac{|D|}{|V|}\rceil$ is not linearly related to the success rate of watermark embedding. In other words, increasing the vocabulary size does not necessarily make watermark embedding easier. In the current setting, $0.5$ appears to be the sweet spot value.
Furthermore, we explore the impact of the maximum number of output tokens of the LLM on watermark embedding. We kept the base setting of $\lceil\frac{|D|}{|V|}\rceil$. Simultaneously, we test the maximum number of output tokens by setting it to $[200,300,400,500,600]$, respectively. Our results are shown in Fig. LABEL:fig:rq1.(b). From the results, we can observe that as the maximum number of output tokens of the LLM increases, the success rate of watermark embedding also tends to increase. However, there is an evident diminishing marginal effect.
Finally, we explore the impact of hash value $H$ on the watermark Embedding Success Rate. In MCGMark, the hash value is continuously varied to ensure the randomness of vocabulary partitioning. Thus, for our purposes, we modify the hash strategy to a fixed approach, setting the hash value to $7,775$ and $666$, respectively. All other settings remained unchanged. Only nine watermarks were successfully embedded under the hash value of $7,775$, yielding an embedding success rate of $18\%$. Conversely, under the hash value of $666$, $19$ watermarks were successfully embedded, resulting in a success rate of $38\%$. We further test scenarios with $H$ values of $15,485,863$ and two, resulting in embedding success rates of $36\%$ and $44\%$, respectively. From this phenomenon, two conclusions can be drawn: (1) The choice of hash value significantly impacts the watermark embedding success rate, providing an opportunity for SSP to enhance performance by selecting appropriate hash values. (2) MCGMark’s pseudo-random hash strategy is effective and can substantially improve the watermark embedding success rate compared to fixed hash values. Utilizing MCGMark’s strategy, SSP no longer needs to expend additional time exploring hash values.
TABLE IV: The impact of fixed hash values on watermark embedding success rate.
@@
6 Limitations and Threats to validity
6.1 Limitations
In Section 5.4, while the watermark demonstrates robust performance in most scenarios, it occasionally fails to defend against certain attacks. Our analysis of these failures reveals that incomplete token adaptation during watermark embedding is a primary issue. Currently, we rely on regular expressions for token recognition, but LLMs often have extensive vocabularies (3.2K tokens or more), making comprehensive adaptation challenging. In future work, we plan to develop a more robust token recognition approach or seek support from SSPs to better address this problem.

### v2

@@ 5. Experiments
  • RQ5. What is the time overhead of MCGMark?
For RQ1, we evaluate the watermark embedding success rate and detection success rate of MCGMark across different LLMs, while comparing it with other state-of-the-art watermarking methods. For RQ2, we compare the code generation quality of MCGMark and its baselines with the CodeBLEU (Ren et al., 2020). Additionally, we conduct a user study to assess the impact of watermark embedding on the quality of generated code. For RQ3, we evaluate the robustness of MCGMark and its baselines under adversarial scenarios. Furthermore, for RQ4, we investigate the influence of key parameters on the embedding success rate of MCGMark and examined the efficiency of watermark generation. Finally, in RQ5, we investigate the impact of MCGMark on time overhead by using different LLMs and comparing against the baselines.
5.1. Experiment Setup
In this part, we present the experimental setup and the baseline methods used for comparison.
LLMs: We evaluate our watermarking strategy on three state-of-the-art LLMs: DeepSeek-Coder-6.7b-instruct (Guo et al., 2024), StarCoder 2-7b-hf (Lozhkov et al., 2024), and CodeLlama-7b-hf (Rozière et al., 2023). These LLMs were selected because they are open-source, which allows us to access their vocabularies and implement MCGMark. Furthermore, they have demonstrated strong performance across multiple benchmarks (Guo et al., 2024; Shin et al., 2024).
@@ 5.1. Experiment Setup
Parameters of MCGMark: We set the number of LLM maximum output token, $L$, to $400$. Additionally, we set $\lceil\frac{|D|}{|V|}\rceil=0.5$. The total length of the watermark, $X$, is set to $24$ bits. And the watermark information is randomly generated. $Q3$ is set to $0.75$, and $Q1$ is set to $0.25$ (Bondarenko et al., 2024).
Evaluation Dataset: We evaluate MCGMark using MCGTest and CodeBLEU (Ren et al., 2020). MCGTest consists of $406$ malicious code prompts, including real instances generated by LLMs and malicious code collected from high-quality open-source repositories. CodeBLEU is an evaluation framework designed for code generation tasks. It integrates traditional n-gram matching, syntax matching, and semantic matching to measure code generation quality comprehensively.
Baselines: We introduce three state-of-the-art LLM watermarking techniques as baselines for comparison: WLLM (Kirchenbauer et al., 2023a), MPAC (Yoo et al., 2024), and PostMark (Chang et al., 2024). WLLM embeds zero-bit watermarks by partitioning the LLM vocabulary into red and green subsets, ensuring the model selects tokens exclusively from the green subset. This approach focuses solely on identifying whether code is generated by an LLM. MPAC extends WLLM on embedding multi-bit watermarks by controlling token selection between vocabularies. Finally, PostMark is a post-processing watermarking method that uses a private vocabulary substitution table to indicate whether the content is LLM-generated. Like WLLM, PostMark is also a zero-bit watermarking scheme.
Parameters of baselines: For the parameter configurations of the three baselines, we follow the default settings as specified in their original works. Specifically, for WLLM, $\gamma$ was set to $0.25$ and $\delta$ to $2$. To ensure fairness, WLLM’s generated token count was limited to $400$, and it utilized the same hash function as MCGMark. For MPAC, while maintaining its default settings, the watermark bit length was set to $24$, and the token count was also limited to $400$.
@@
Implementation: We implement MCGMark in Python. All experiments are conducted on a workstation with $128$ CPU cores and $8$$\times$ NVIDIA A800 (80G) GPUs.
5.2. RQ1: Effectiveness of MCGMark
To answer RQ1, we first evaluate the watermark embedding success rate and detection success rate of MCGMark across different LLMs. This evaluation not only assesses the effectiveness of watermark embedding in LLMs but also demonstrates the adaptability of MCGMark to various models. Additionally, we compare the performance of different watermarking schemes under identical settings to further highlight the capabilities of MCGMark.
Watermark Embedding Success Rate. In this part, we apply MCGMark to different LLMs and tested the success rate of watermark embedding using the MCGTest dataset. Watermark embedding is attempted three times. As shown in Table 2, the average watermark embedding success rate across the three LLMs was 85.2%. Specifically, MCGMark achieved the highest embedding success rate with DeepSeek-Coder at 88.9%, while the success rates for CodeLlama and StarCoder-2 were 79.6% and 87.2%, respectively. These results demonstrate that MCGMark is adaptable to different LLMs, aligning with the theoretical findings reported by Kirchenbauer et al (Kirchenbauer et al., 2023b). Furthermore, the results validate the effectiveness of MCGMark in watermark embedding.
@@ 5.2. RQ1: Effectiveness of MCGMark
Table 2. Watermark Embedding Success Rate of MCGMark with various LLM.
LLM Model Sucess Rate.(%)
Deepseek-Coder 6.7b-instruct 88.9
StarCoder-2 7b-hf 87.2
CodeLlama 7b-hf 79.6
Average / 85.2
@@
Subsequently, to further explore the effectiveness of MCGMark, we analyze the $45$ instances where MCGMark failed to embed watermarks on DeepSeek-Coder. We find that $17$ tasks failed due to the generated code being too short to embed the watermark. Additionally, $25$ tasks failed because the generated code triggered numerous traits, resulting in a high number of assignment statements and comments, which hindered watermark embedding. Another three prompts reject the malicious code generation request, resulting in empty content. Higher folding watermark encoding could potentially solve the issue of shortcode generation failures, though this is not the focus of this paper. The failures related to code structure primarily stemmed from our restriction of LLMs to generate a maximum of $400$ tokens. In practical applications, LLMs typically generate $2048$ to $4096$ tokens or even more (e.g., DeepSeek-Coder-6.7b can generate up to 64K (Guo et al., 2024)). We retest the $25$ tasks that fail due to code structure with a maximum length setting of $2048$ tokens and successfully embed watermarks in $21$ of them, achieving an embedding success rate of $84\%$. Therefore, under conditions where token numbers are unrestricted, MCGMark’s watermark embedding success rate could reach approximately $94.1\%$ across $406$ tasks.
Watermark Detecting Success Rate. In this part, we evaluate the detection success rate of MCGMark and compared it with other watermarking strategies. To ensure fairness, we analyze the performance of the DeepSeek-Coder-6.7b-instruct model with different watermarking strategies on the MCGTest dataset. The results are presented in Table 3. As shown, MCGMark achieved the highest watermark detection success rate. Both WLLM and MPAC also demonstrated relatively high detection success rates. However, the detection success rate of PostMark, a post-processing watermarking method, was significantly lower. This can be attributed to PostMark’s reliance on the substitution model’s capability and the adaptability of its maintained substitution table.
Table 3. Watermark detecting Success Rate of various watermark.
Watermark In-process Multibit Code-aware Detect Rate.(%)
[…]
@@
Analysis of Detection Failures. In this part, we first analyze the scenarios where MCGMark failed in watermark detection. Subsequently, we examine the impact of tolerance bits on the watermark detection success rate. For the first study, Out of the $361$ tasks where MCGMark successfully embed watermarks, $353$ watermarks are detected successfully, resulting in a detection success rate of approximately $97.8\%$. In contrast, WLLM does not support false positive rate checking. The eight instances where MCGMark failed to detect watermarks can be attributed to discrepancies in token splitting by the tokenizer. This issue leads to errors when verifying tokens against the vocabulary, returning incorrect results. This limitation is inherent to SSP, and we cannot improve the detection success rate by modifying MCGMark.
For the second part, we conduct a detailed evaluation of the impact of tolerance bits on watermark detection. Specifically, we assess the detection success rate of watermarks without applying tolerance bits in 361 cases where watermark embedding was successful. We observe that without error-correcting bits, the watermark is rarely detected successfully. This outcome can be attributed to two main reasons. First, the inherent randomness in vocabulary partitioning makes it difficult to ensure that outlier tokens consistently fall into the vocabulary group aligned with the watermark encoding. Second, the total watermark length is 24 bits, and under this setting, the theoretical probability of successful detection without error correction is only $0.19\%$. These results further highlight the necessity of incorporating error-correcting bits in our watermark design.
Answer to RQ 1: MCGMark achieved a watermark embedding success rate of over 85% and a watermark detection success rate of 97.8%. Compared to other watermarking methods, it demonstrated superior performance. Moreover, MCGMark is model-agnostic and does not rely on specific LLMs.
5.3. RQ2: Impact on LLM Code Generation
To address RQ2, we first evaluate the quality of code generated under different watermarks with CodeBLEU (Ren et al., 2020), a widely adopted framework for assessing code quality. Additionally, we conduct a user study to further investigate the impact of MCGMark on the quality of code generation.
CodeBLEU Result. We assess how different watermarking strategies affect code quality using CodeBLEU, comparing results to code generated without watermarking. Inspired by previous work (Zhang and Koushanfar, 2024; Guan et al., 2024), we evaluate the impact of different watermarking methods on code quality by comparing CodeBLEU scores before and after watermark embedding. In this evaluation, a higher CodeBLEU score indicates a smaller impact of the watermark on the LLM. To ensure fairness, we address a specific behavior of PostMark. Since PostMark sometimes converts code entirely into plain text during watermark embedding—an inherent characteristic of its design—we assign a score of 0 to such cases. For all successfully watermarked code segments, we compute CodeBLEU scores before and after embedding and report the average score. The results are presented in Table 4.
In Table 4, the second column indicates whether the watermark is multi-bit. The third column shows whether the watermark embedding or detection process depends on external databases or requires loading additional models. As shown in Table 4, all watermarking strategies, including MCGMark, result in a decline in the code generation quality of LLMs. MCGMark achieved higher scores, indicating a smaller impact of the watermark on code generation. Furthermore, compared to other watermarking strategies, MCGMark supports embedding a greater amount of multi-bit information while having a less pronounced impact on code quality.
Table 4. CodeBLEU results of various watermark.
Watermark Multibit(digits) No external dependencies Core
[…]
Additionally, some watermarking strategies, such as PostMark, rely on external databases or models, requiring additional resources to be loaded during watermark embedding and detection. In contrast, MCGMark operates independently of such external dependencies.
@@ 5.3. RQ2: Impact on LLM Code Generation
User Study. We further conduct a user study to evaluate the impact of MCGMark on the quality of LLM-generated code. We primarily assess the code generation quality of the LLM to explore the impact of MCGMark on the code generation quality. In this part, we randomly select $50$ tasks from the $346$ successfully embedded tasks of MCGTest. Furthermore, we obtain $100$ code segments, with $50$ generated by the model using our watermarking strategy and $50$ without. We invite $10$ developers with at least $4$ years of development experience (excluding co-authors), including $6$ Ph.D. students, $2$ undergraduate students, and $2$ software engineers specializing in computer-related fields, to participate in our evaluation. We randomly shuffle the order of the $50$ code pairs and further shuffle the code order within each pair. We ask the developers to identify the code they believe contains a watermark for each code pair. We collect a total of $1000$ valid responses and organize the accuracy of these $1000$ responses, as shown in Table 5.
Table 5. The results of the user study on distinguishing watermark and unwatermarked code.
participants 1 2 3 4
Correct Number / Success Rate 22/44% 27/54% 21/42% 26/52%
5 6 7 8 9 10 Avg
30/60% 19/38% 21/42% 24/48% 21/42% 28/56% 23.9/47.8%
@@
The recognition accuracy of $10$ participants for $500$ pairs of watermark/unwatermark code, totaling $1000$ segments, is $47.8\%$, which is close to random sampling. Moreover, the independent recognition rates of the $10$ participants, excluding the highest value (participant 5) and the lowest value (participant 6), fluctuate between $40\%$ to $50\%$. Furthermore, all these rates fall within the $95\%$ confidence interval ($[21.972,27.828]$). So, we can conclude that experienced practitioners with long-term development experience cannot correctly distinguish between watermark and unwatermark codes. This also demonstrates the stealthiness of our watermark and further confirms that the impact of MCGMark on code quality can be considered negligible.
@@
Answer to RQ 2: MCGMark preserves higher code generation quality compared to baseline methods while embedding watermarks, without noticeably impacting the normal functionality of the LLM.
5.4. RQ3: Resistance to Tampering
To address RQ3, we first evaluate the robustness of watermarks against eight types of attacks. Subsequently, we conduct a detailed analysis of the results for MCGMark.
Robustness Test. In this part, our main focus is on the robustness of the watermark in Section 4.4. Based on the literature (Li et al., 2022), we primarily consider the following two types of attacks, as shown in Table 6. These $8$ attack types cover the majority of modifications that typically employ against code watermarks (Haq and Caballero, 2021; Min and Li Ping, 2019). These eight types of attacks do not compromise the functionality of the code, making them more representative of the behavior of malicious developers with low coding proficiency. We evaluate the performance of $50$ successfully watermarked in MCGTest codes under $8$ attacks, particularly whether these attacks affect the embedded watermark elements. For each attack, we conduct three attack instances. We carry out a total of $1200$ attacks.
Table 6. Types and descriptions of attacks.
@@ 5.4. RQ3: Resistance to Tampering
(2) modify/remove inputs and outputs Modifying or removing the content of program inputs and outputs.
(3) modify/remove comments Modifying or removing single-line comments or multi-line comments.
@@
We conduct the aforementioned attacks on four watermarking strategies, resulting in a total of 4,800 attacks. The results are shown in Table 7. As observed, MCGMark achieved relatively favorable results across all eight types of attacks, whereas other watermarking strategies often exhibited weaker robustness under specific types of attacks. For instance, WLLM demonstrated poor robustness when faced with modifications involving longer text, such as removing or adding comments. This is because WLLM relies on statistical rules to determine whether a piece of code contains a watermark. Large-scale removal of comments, while not affecting code functionality, significantly impacts WLLM’s watermark detection performance.
Similarly, PostMark shows weaker robustness against attacks that modify or remove code elements. MPAC performs more robustly across the eight attack types, likely due to mechanisms like List Decoding, but struggles with Type 1 attacks. Overall, MCGMark demonstrates consistently high robustness.
Robustness Analysis. MCGMark failed to defend against attacks, to further investigate its robustness. In $1200$ attacks, we achieve a complete defense against Attacks 3, 6, 7, and 8. However, Attacks 4 and 1 failed $13$ times and $9$ times, respectively, resulting in defense success rates of $91.3\%$ and $94\%$. Upon carefully examining the failed instances, we identify a flaw in our token matching strategy during the implementation of Pattern 1-7. We rely on a generic regular expression matching approach, which proves inadequate for matching the tokens generated by LLMs due to their deviation from human language rules. The same phenomenon also affects the defense against Attack 2, resulting in an accuracy rate of $93.3\%$. Fortunately, this issue can be resolved by designing a more powerful matching scheme or SSP adapting the token vocabulary for the watermark.
@@
Table 7. Robustness on various watermark.
Attack WLLM PostMark MPAC MCGMark
Type 1.(1) 150/150 148/150 143/150 142/150
Type 1.(2) 150/150 143/150 122/150 139/150
Type 1.(3) 146/150 105/150 114/150 150/150
Type 1.(4) 141/150 133/150 126/150 137/150
Type 1.(5) 149/150 137/150 124/150 129/150
Type 2.(6) 111/150 74/150 142/150 150/150
Type 2.(7) 116/150 118/150 150/150 150/150
Type 2.(8) 147/150 113/150 141/150 150/150
Total 1110/1200 971/1200 1062/1200 1147/1200
Additionally, it is worth emphasizing that malicious developers may attempt to modify the code generated by LLMs. However, as long as the LLM use our watermark, the watermark information cannot be removed from the code.
@@ 5.5. RQ4: Impact of Settings
Hyperparameters. We examine the impact of three hyperparameters on watermark embedding success. We randomly select 50 prompts from the 406 tasks and vary one hyperparameter at a time, keeping others fixed. First, we analyze $\lceil\frac{|D|}{|V|}\rceil$, which controls the proportion of vocabulary $\mathbb{A}$ or $\mathbb{B}$ during partitioning. Following (Kirchenbauer et al., 2023a), we fix the LLM’s output length at 400 and set $\lceil\frac{|D|}{|V|}\rceil$ to $[0.25,0.375,0.5,0.625,0.75]$. As shown in Fig. LABEL:fig:rq1(a), indicate that the success rate does not increase monotonically with the vocabulary ratio. In other words, a larger vocabulary does not necessarily facilitate watermark embedding. Under the current configuration, a ratio of $0.5$ yields the best performance.
Maximum number of output tokens. Furthermore, we explore the impact of the maximum number of output tokens of the LLM on MCGMark. We keep the base setting of $\lceil\frac{|D|}{|V|}\rceil$. Simultaneously, we test the maximum number of output tokens by setting it to $[200,300,400,500,600]$. Our results are shown in Fig. LABEL:fig:rq1.(b). From the results, we can observe that as the maximum number of output tokens of the LLM increases, the success rate of watermark embedding also tends to increase. However, there is an evident diminishing marginal effect.
Hash value. Finally, we explore the impact of hash value $H$ on the watermark Embedding Success Rate. In MCGMark, $H$ is continuously varied to ensure the randomness of vocabulary partitioning. For comparison, we adopt a fixed-hash strategy using values $7,775$ and $666$, keeping all other settings unchanged. Only nine watermarks were successfully embedded under the hash value of $7,775$, yielding an embedding success rate of $18\%$. Conversely, under $H$ on $666$, $19$ watermarks were successfully embedded, resulting in a success rate of $38\%$. We further test scenarios with $H$ values of $15,485,863$ and two, resulting in embedding success rates of $36\%$ and $44\%$, respectively. This phenomenon leads to two key conclusions: (1) The choice of hash values affects the success rate of watermark embedding. Since hash values are typically private to the SSP, selecting an appropriate initial hash value can further improve watermark embedding performance. (2) MCGMark’s pseudo-random hash strategy proves highly effective, significantly improving the embedding success rate compared to fixed hash values. By adopting MCGMark’s approach, SSP can eliminate the need to spend additional time searching for optimal hash values.
Table 8. The impact of fixed hash values on watermark embedding success rate.
@@
5.6. RQ5: Extra time overhead
To address RQ5, we analyze the additional time overhead introduced by MCGMark. Specifically, we examine the additional time overhead of MCGMark when applied to different LLMs and compared it with other watermarking strategies when using the same LLM.
Overhead with various LLM. We evaluate the time overhead introduced by MCGMark on three state-of-the-art LLMs. To clearly illustrate the impact of MCGMark, we also measure the time overhead of the LLMs without watermarking. The evaluation was conducted on all 406 prompts from MCGTest, and the average results were calculated to provide a more intuitive comparison. The results are shown in Table 9. From the results, we can observe that MCGMark does introduce additional time overhead, and this overhead shows some correlation with the characteristics of the respective LLMs.
Table 9. Time overhead of MCGMark across different LLMs. (wo) indicates the setting without watermarking, and (w) indicates the setting with watermarking.
LLM Overhead (wo) Overhead (w) Multiple
Deepseek-Coder 14.04 103.48 7.37
StarCoder-2 11.69 120.55 10.31
CodeLlama 9.65 91.31 9.87
Overhead of various watermark. We further evaluate the time overhead by different watermarking strategies, as shown in Table 10. To better illustrate the result, we measure both the embedding overhead and detection overhead for each watermarking strategy. From the results, we observe that all watermarking strategies introduce additional time overhead. Among them, WLLM, the online zero-bit watermarking method, is the most lightweight. In contrast, PostMark incurs higher time overhead as it requires invoking the LLM during both code generation and watermark embedding. The overall time overhead of MPAC and MCGMark is similar, with both methods introducing more latency during watermark embedding. Meanwhile, MPAC requires additional operations in its implementation pipelines, leading to higher embedding delay. Overall, MCGMark remains competitive compared to other baselines.
Table 10. Time overhead of various watermarking strategies. (E) indicates the time required for watermark embedding, (D) indicates the time for watermark detection, and (T) represents the total overhead.
Watermark Overhead (E) Overhead (D) Overhead (T)
No watermark / / 14.04
WLLM 41.62 1.29 42.91
PostMark 202.48 0.21 202.69
MPAC 112.31 2.14 114.45
MCGMark 99.12 4.36 103.48
Answer to RQ 5: MCGMark introduces a measurable level of additional time overhead. However, it still achieves competitive results compared to the baselines.
6. Limitations
Defense Against Specific Attacks. In Section 5.4, the experiments demonstrate that the watermark exhibits overall robustness and can withstand most tampering attacks, but it may fail in certain specific cases. Analysis shows incomplete token adaptation due to reliance on regex-based recognition of extensive LLM vocabularies (over 3.2K tokens) as a primary issue. Improving token adaptation through refined recognition techniques and SSP-based support constitutes a direction for future research. Additionally, inserting non-watermarked human-written code can contaminate detection, though this typically changes the code’s functionality, making it effectively non-LLM-generated. Enhanced watermark truncation and extraction techniques are considered promising directions for future studies to address such attacks.